OpenStudy (anonymous):

the x and y-intercepts of f(x)=sin(x+π/2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I mean how would you find intercepts normally?

OpenStudy (tkhunny):

y-intercept, if any, substitute x = 0. That's it. With a function, this is not normally very difficult. x-intercept, if any, substitute y = 0. This can be a bit trickier.

OpenStudy (anonymous):

so the x-intercept would be (π/2,0)?

OpenStudy (tkhunny):

No. That might be one of them. \(\sin(x + \pi/2) = 0 \implies x + \pi/2 = 0 + k\pi\) where k is ANY Integer. Thus, \(x = -\pi/2 + k\pi\) and since we can pick ANY integer, we see that there are infinitely many x-interecepts.

OpenStudy (anonymous):

ok thank you!